OpenStudy (anonymous):

Evaluate \[\int\limits_{e}^{4e} 1/x dx\]

OpenStudy (richyw):

do you know the antiderivative of 1/x?

OpenStudy (anonymous):

I got this one already \[\ln \left| 4e \right|- \ln \left| e \right|\] but then I got stuck

OpenStudy (richyw):

use your log rules

OpenStudy (richyw):

\[\log a - \log b = \log\frac{a}{b}\]

OpenStudy (anonymous):

I did not learn about it yet

OpenStudy (richyw):

hmm. you should have learned about those before you did calculus. you probably will want to refresh on those. but I gave the formula you need for this question anyways.

OpenStudy (anonymous):

we can use ln rule then

OpenStudy (richyw):

oh yeah, I wrote log instead of ln (although it doesn't matter for the rule).

OpenStudy (anonymous):

ok...I got it

OpenStudy (richyw):

it works for any logs as long as the base is the same

OpenStudy (anonymous):

I'm turning crazy when I took like the whole week for cal..

OpenStudy (anonymous):

thank you so much

OpenStudy (richyw):

no worries.

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

\(\log(4e)-\log(e)=\log(4)+\log(e)-\log(e)=\log(4)\)

OpenStudy (kc_kennylau):

Other log rules for you: \[\Large\log_na+\log_nb=\log_nab\]\[\Large\frac{\log_na}{\log_nb}=\log_ba\]

OpenStudy (tkhunny):

Super good try on \(\log|4e| - \log|e|\). If we knew nothing of "e", this would be exactly correct. Since we know \(e > 0\), the Absolute Values are not necessary.

OpenStudy (tkhunny):

You should not be in Calculus without a background in logarithms. Something seriously wrong, there. Find an algebra book - maybe you saved one from a previous class. There will be a section in there on logarithms. Why someone decided to skip the section on logarithms is a grave mystery. Maybe it was assumed that you never would get to the calculus? {shakes head}

OpenStudy (anonymous):

Thank you so much tkhunny, I'm planning to review all my math acknowledge on the summer.....